# Stochastic Online Metric Matching: Adversarial is no Harder than Stochastic

### v1

@@ 4 Euclidean Space with Smooth Distributions
Next, we consider balanced markets with smooth distributions in Section 4.2. Recall that when $d=1$, an $O(1)$-competitive algorithm in balanced markets for general distributions [^5] In Section C, we also give an alternative $O(1)$-competitive algorithm for $d=1$ in balanced markets with general distributions. is given by [GGPW19]. Moreover, the above algorithm admits an incentive ratio of $O((\log\log\log n)^{2})$ for general metrics and distributions. However, for $d=2$ and smooth distributions, we are only able to estimate the optimal offline solution up to a $\Theta(\sqrt{\log n})$ factor, which prevents our approach from achieving an $o(\sqrt{\log n})$ competitive ratio. Hence, we only focus on the case where $d\geq 3$. In particular, we present an $O(1)$-competitive algorithm for all $d\geq 3$.
@@
4.1 Warm-up: Balanced Markets with Uniform Distribution
In this subsection, we consider balanced markets and assume $\mathbb{D}$ to be uniform distribution. We show that combining Theorem 1 with the Hierarchical Greedy algorithm of [Kan22] results in an $O(1)$-competitive algorithm for all $d\geq 1$.
We present our algorithm in the following theorem.
@@ 4.1 Warm-up: Balanced Markets with Uniform Distribution
Suppose that $m=n$. For all $d\geq 1$, an $O(1)$-competitive algorithm exists for adversarial servers and uniformly distributed requests.
@@
Recall that [Kan22] gives an algorithm with a competitive ratio of $O(1)$ for all $d\geq 1$ when both servers and requests follow uniform distribution. Combining the above algorithm with Theorem 1 concludes the proof. ∎
@@ 4.2 Balanced Markets
In this subsection, we assume $d\geq 3$ and consider balanced markets with smooth distributions. We will prove that the $\mathsf{SOAR}$ algorithm admits an $O(1)$ competitive ratio for $d\geq 3$ when servers and requests are randomly drawn from $\mathbb{D}$. Together with Theorem 1, this leads to an $O(1)$-competitive algorithm for $d\geq 3$ with adversarial servers and stochastic requests.
@@
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are randomly drawn from $\mathbb{D}$, an $O(1)$-competitive algorithm exists for $d\geq 3$.
The rest of this section is devoted to proving Theorem 3. By Theorem 1, it suffices to show that the $\mathsf{SOAR}$ algorithm is $O(1)$-competitive for $d\geq 3$ when both servers and requests are randomly drawn from $\mathbb{D}$. We compute the expected cost of the $\mathsf{SOAR}$ algorithm in the following lemma. Even though we only consider $d\geq 3$ in this subsection, the costs of the $\mathsf{SOAR}$ algorithm for $d\leq 2$ computed by Lemma 7 will be useful afterward.
@@ 4.3 Unbalanced Markets
Suppose that $n\leq m\leq O(n)$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are randomly drawn from $\mathbb{D}$, an algorithm exists with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
@@
By Lemma 7, an algorithm ${\mathcal{A}}$ exists such that
Moreover, by Theorem 1, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $n\leq|S|\leq O(n)$. Combining the lower bounds for $\mathrm{OPT}(S,n)$ given by Lemma 4 concludes the proof. ∎
5 Euclidean Space with General Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance. We do not make any assumptions on the distribution and derive algorithms with nearly optimal regret bounds for all $d\geq 1$ in unbalanced markets with at most linearly many excess servers. In particular, our results recover the nearly optimal regret bounds by [CKKZ23], which is established in balanced markets with servers and requests following different distributions. In the weaker setting studied by [CKKZ23], the regret bound is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$, implying that our regret bounds are also nearly optimal.
@@ 5 Euclidean Space with General Distributions
Suppose that $n\leq m\leq O(n)$ and $\mathbb{D}$ is an arbitrary distribution on $[0,1]^{d}$. An algorithm ${\mathcal{A}}^{\prime}$ exists that, for every set of servers $S$ with $|S|=m$, admits a regret of
@@
By Theorem 1, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $n\leq|S|\leq O(n)$. As a result, the regret of ${\mathcal{A}}^{\prime}$ equals
@@ 6.1 Applications on Hierarchically Separated Tree Metrics
Proof.
Therefore,
which concludes that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing. ∎
The following corollary is a direct consequence of combining Theorem 6 and Lemma 9 with the competitive ratio guarantee of the MNP algorithm in balanced markets [MNP06].
@@
Proof.
concluding that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing. ∎
Combining Theorem 6 and Lemma 10 with the competitive ratio guarantee of the Random Subtree algorithm in balanced markets [GL12], we derive the competitive ratio guarantee of the Random Subtree algorithm in unbalanced markets in the following corollary.
@@ 7 Discussion and Future Directions
In this paper, for balanced markets, we achieve an $O(1)$ competitive ratio for the Euclidean space with smooth distributions and $d\neq 2$. Nevertheless, it still remains open to give an $O(1)$-competitive algorithm for general metrics and distributions. The potential approaches for achieving this goal comprise (1) improving the analysis of the algorithm of [GGPW19], which is proved to be $O((\log\log\log n)^{2})$-competitive without any super-constant lower bound known, and (2) applying Theorem 1 and the $\mathsf{SOAR}$ algorithm to the simplified model given by [GGPW19] (Lemma 15). The main technical barrier underlying the latter approach lies in the difficulty of characterizing $\mathrm{OPT}(n)$ for general metrics and distributions.
@@
We also derive an algorithm in unbalanced markets for the Euclidean space and smooth distributions with improved competitive ratios. Our algorithm admits an optimal constant competitive ratio for $d\geq 3$, leaving the improvement for $d\leq 2$ an intriguing future direction. Given that the line metric, i.e., $d=1$, is typically the most tractable case [GGPW19, BFP23], it is surprising that it becomes the most intractable one with the presence of excess servers, rendering its resolution more appealing.
@@
Moreover, for adversarial servers and requests, we show that the competitive ratio of every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets is preserved in unbalanced markets. However, our characterization fails to capture all the existing algorithms. In this most extensively studied setting, our understanding of the achievable guarantees in unbalanced markets remains very limited, and nothing beyond our results is known. An exciting future direction is to generalize the performance of the existing algorithms with provable guarantees in balanced markets to unbalanced markets.
Finally, we showcase the generality of our reduction by establishing positive results for cost functions only approximately satisfying the triangle inequality in Appendix B. Conversely, a strong impossibility result for achieving a bounded competitive ratio without any assumption on cost functions exists [GGPW19]. Our results can be viewed as pushing the boundary within which a bounded competitive ratio is achievable, and it would be interesting to further characterize the boundary between possibility and impossibility.
Acknowledgments
We thank Yash Kanoria, Akshit Kumar, Tristan Pollner, Shayan Talaei, and Yilun Chen for their helpful discussions, and Kiana Ashgri for pointing out the proof of Lemma 13.
@@ A.3 Proof of Lemma 4
Appendix B Cost Functions Beyond Triangle Inequality
In previous sections, we consider metric space where the cost function satisfies the exact triangle inequality. In this section, we showcase the generality of our reduction in Section 3 by proving positive results when the cost function only approximately satisfies the triangle inequality. Specifically, we say that a cost function $\delta$ on $X$ satisfies the $\eta$-approximate triangle inequality for $\eta>0$ if for all $x,y,z\in X$,

### v2

@@ 4 Stochastic Setting: Euclidean Space with Smooth Distributions
Next, we consider balanced markets with smooth distributions in Section 4.2. Recall that when $d=1$, an $O(1)$-competitive algorithm in balanced markets for general distributions is given by [GGPW19]. Moreover, this algorithm is $O((\log\log\log n)^{2})$-competitive for general metrics and distributions. However, for $d=2$ and smooth distributions, we are only able to estimate the optimal offline cost up to a $\Theta(\sqrt{\log n})$ factor, which prevents our approach from achieving an $o(\sqrt{\log n})$ competitive ratio. Hence, we only focus on the case with $d\geq 3$, where we present an $O(1)$-competitive algorithm.
@@
4.1 Prelude: Balanced Markets with Uniform Distribution
In this subsection, we consider balanced markets and assume $\mathbb{D}$ to be uniform distribution. In the following theorem, we show that combining Theorem 2 with the Hierarchical Greedy algorithm of [Kan22] results in an $O(1)$-competitive algorithm for all $d\geq 1$.
@@ 4.1 Prelude: Balanced Markets with Uniform Distribution
Suppose that $m=n$. For $d\geq 1$, an $O(1)$-competitive algorithm exists for adversarial servers and uniformly distributed requests.
@@
Recall that [Kan22] gives an $O(1)$-competitive algorithm for $d\geq 1$ when both servers and requests follow uniform distribution. Combining the above algorithm with Theorem 2 concludes the proof. ∎
@@ 4.2 Balanced Markets
In this subsection, we assume $d\geq 3$ and consider balanced markets with smooth distributions. We will prove that the SOAR algorithm is $O(1)$-competitive for $d\geq 3$ when servers and requests are randomly drawn from $\mathbb{D}$. Together with Theorem 2, this leads to an $O(1)$-competitive algorithm for $d\geq 3$ with adversarial servers and stochastic requests.
@@
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$, an $O(1)$-competitive algorithm exists for $d\geq 3$.
The rest of this subsection is devoted to proving Theorem 5. By Theorem 2, it suffices to show that the SOAR algorithm is $O(1)$-competitive for $d\geq 3$ when both servers and requests are randomly drawn from $\mathbb{D}$. We compute the expected cost of the SOAR algorithm in the following lemma. Even though we only consider $d\geq 3$ in this subsection, the costs of the SOAR algorithm for $d\leq 2$ computed by Lemma 8 will be useful afterward.
@@ 4.3 Unbalanced Markets
Suppose that $n\leq m\leq O(n)$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$, an algorithm exists with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
@@
Let ${\mathcal{A}}$ be the algorithm as per Lemma 8. By Theorem 2, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $n\leq|S|\leq O(n)$. Combining the lower bounds for $\mathrm{OPT}(S,n)$ given by Lemma 4 concludes the proof. ∎
5 Stochastic Setting: Euclidean Space with General Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance. We do not make any assumptions on the distribution and achieve nearly optimal regret bounds for all $d\geq 1$ in unbalanced markets with an arbitrary number of excess servers. In particular, our results recover the nearly optimal regret bounds by [CKKZ23], which is established in balanced markets with servers and requests following different distributions. In the weaker setting studied by [CKKZ23], the regret bound is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$, implying that our regret bounds are also nearly optimal.
@@ 5 Stochastic Setting: Euclidean Space with General Distributions
Suppose that $n\leq m$ and $\mathbb{D}$ is an arbitrary distribution over $[0,1]^{d}$. An algorithm ${\mathcal{A}}^{\prime}$ exists that, for every set of servers $S$ with $|S|=m$, admits a regret of
@@
By Theorem 2, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. As a result, the regret of ${\mathcal{A}}^{\prime}$ equals
@@
6 Prediction Setting
In this section, we adopt the same setting and notations as in Section 3.2. We will present an algorithm with desirable consistency and robustness guarantees, whose performance also smoothly depends on the prediction error. Recall that $\eta$ denotes the prediction error and $\mathrm{OPT}$ denotes the optimal offline cost, and our main result in this section is presented as follows.
Theorem 8.
Assume that $m=n$. Given an algorithm ${\mathcal{A}}$ that does not use predictions and is $\alpha$-competitive for a particular metric, there exists an algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions and is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive for the same metric. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
The proof of Theorem 8 is deferred to Section 6.1, and we now apply Theorem 8 to the existing algorithms in the literature. Recall that without predictions, there exist deterministic $(2n-1)$-competitive algorithms for general metrics [KP93, KMV94, Rag16], a deterministic $O(\log n)$-competitive algorithm for the line metric [Rag18], a randomized $O(\log^{2}n)$-competitive algorithm for general metrics [BBGN14], and a randomized $O(\log n)$-competitive algorithm for doubling metrics [GL12]. As a corollary, we derive the following results.
Corollary 1.
For general metrics, there exist a deterministic $O(\min\{n,1+\frac{\eta\cdot n}{\mathrm{OPT}})\}$-competitive algorithm and a randomized $O(\min\{\log^{2}n,1+\frac{\eta\cdot\log^{2}n}{\mathrm{OPT}}\})$-competitive algorithm. For the line metric, there exists a deterministic $O(\min\{\log n,1+\frac{\eta\cdot\log n}{\mathrm{OPT}})\}$-competitive algorithm. For doubling metrics, there exists a randomized $O(\min\{\log n,1+\frac{\eta\cdot\log n}{\mathrm{OPT}})\}$-competitive algorithm.
6.1 Proof of Theorem 8
Recall that we have given a non-robust algorithm with the desired consistency guarantee in Theorem 3, and now we show how to robustify it, which leads to the desired algorithm.
A powerful machinery for deterministically combining several algorithms into one in a black-box manner is proposed by [FRR94] for the $k$-server problem and is subsequently applied by [LV21, ACE+23b] under the context of online algorithm with predictions. We will use the same approach to combine our consistent but non-robust algorithm with the given algorithm ${\mathcal{A}}$ that does not use predictions, which is by definition robust but inconsistent. Consequently, robustness and consistency are simultaneously achieved.
We adapted the framework for combining several online algorithms generalized by [ACE+23b] to online metric matching, which is a direct consequence of combining [ACE+23b, Theorem 18] and [ACE+23b, Lemma 21].
Lemma 9 (Theorem $18$ and Lemma $21$ in [ACE+23b]).
Given $k$ deterministic algorithms ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$ for online metric matching, for every $1<\gamma\leq 2$, there exists a deterministic algorithm ${\mathcal{A}}^{\mathrm{com}}$ for online metric matching such that for all $S$ and $R$ with $|S|=|R|$,
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,.$
We generalize Lemma 9 to combine randomized online algorithms in the following lemma.
Lemma 10.
Given $k$ randomized algorithms ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$ for online metric matching, for every $1<\gamma\leq 2$, there exists a randomized algorithm ${\mathcal{A}}^{\mathrm{com}}$ for online metric matching such that for all $S$ and $R$ with $|S|=|R|$,
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,.$
Proof.
For each ${\mathcal{A}}_{i}$, recall that ${\mathcal{A}}_{i}$ is randomized, and we use ${\mathcal{A}}_{i}(S,R\mid B)$ to denote the cost incurred by ${\mathcal{A}}_{i}$ with servers $S$ and requests $R$ conditional on the random bits used by ${\mathcal{A}}_{i}$ being $B=b_{1}b_{2}\cdots$. To construct our desired algorithm ${\mathcal{A}}^{\mathrm{com}}$, we first independently sample the random bits $B=b_{1}b_{2}\cdots$ used by ${\mathcal{A}}_{1},\ldots,A_{k}$ and then apply the combination scheme of Lemma 9 to combine ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$. By the guarantee of Lemma 9, for all $S$ and $R$, it holds that
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)$ $\leq\mathbb{E}_{B}\left[\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R\mid B)\}\right]$
$\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathbb{E}_{B}[\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R\mid B)]\}$
$=\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,,$
where the second inequality follows from Jensen’s inequality. ∎
Now, we are ready to prove Theorem 8. We only prove the statement for deterministic algorithms by using Theorem 3 and Lemma 9, and the statement for randomized algorithms can be established analogously by applying Theorem 3 and Lemma 10. Given a deterministic $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions for a particular metric, by Theorem 3, there exists a deterministic $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions for the same metric. By applying Lemma 9 with $\gamma=2$ to combine ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, we obtain a deterministic algorithm ${\mathcal{A}}^{\mathrm{com}}$ incurring a cost of $O(\min\{\mathrm{cost}_{{\mathcal{A}}}(S,R),\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\})$ for all $S$ and $R$ with $|S|=|R|$, which implies that ${\mathcal{A}}^{\mathrm{com}}$ is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive by the competitive ratios of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, concluding the proof.
@@ 7.1 Applications on Hierarchically Separated Tree Metrics
The proof of Lemma 12 is deferred to Section B.1. The following corollary is a direct consequence of combining Theorem 9 and Lemma 12 with the competitive ratio guarantee of the MNP algorithm in balanced markets [MNP06].
@@
The proof of Lemma 13 is deferred to Section B.2. Combining Theorem 9 and Lemma 13 with the competitive ratio guarantee of the Random Subtree algorithm in balanced markets [GL12], we derive the competitive ratio guarantee of the Random Subtree algorithm in unbalanced markets in the following corollary.
@@ 8 Discussion and Future Directions
In this paper, we present a generic approach for online metric matching to utilize the provided distributional knowledge or predictions of request locations, leading to new results in both settings.
@@
When requests are stochastic, for balanced markets, we achieve an $O(1)$ competitive ratio for the Euclidean space with smooth distributions and $d\neq 2$. Nevertheless, it still remains open to give an $O(1)$-competitive algorithm for general metrics and distributions. The potential approaches for achieving this goal comprise (1) improving the analysis of the algorithm of [GGPW19], which is proved to be $O((\log\log\log n)^{2})$-competitive without any super-constant lower bound known, and (2) applying Theorem 2 and the SOAR algorithm to the simplified model given by [GGPW19]. The main technical barrier underlying the latter approach lies in the difficulty of characterizing $\mathrm{OPT}(n)$ for general metrics and distributions. In Appendix D, we provide further evidence that supports the plausibility of this approach by applying it to achieve a constant competitive ratio with tree metrics and general distributions, recovering a result achieved by the algorithm of [GGPW19].
@@
We also derive an algorithm in unbalanced markets with improved competitive ratios when requests are drawn from smooth distributions over the Euclidean space. Our algorithm admits an optimal constant competitive ratio for $d\geq 3$, leaving the improvement for $d\leq 2$ an intriguing future direction. Given that the line metric, i.e., $d=1$, is typically the most tractable case [GGPW19, BFP23], it is surprising that it becomes the most intractable one with the presence of excess servers, rendering its resolution more appealing.
@@
Moreover, for adversarial servers and requests without predictions, we show that the competitive ratio of every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets is preserved in unbalanced markets. However, our characterization fails to capture all the existing algorithms. In this most extensively studied setting, our understanding of the achievable guarantees in unbalanced markets remains very limited, and nothing beyond our results is known. An exciting future direction is to generalize the performance of the existing algorithms with provable guarantees in balanced markets to unbalanced markets.
Several open questions arise from the modeling assumptions in this paper. First, our results assume advance knowledge of $n$, raising the question of whether competitive guarantees are achievable without this information. Next, while we define unbalanced markets as $n\leq m\leq O(n)$ for our competitive ratio results in the stochastic setting, a definition adopted by all prior work, it remains unclear whether these results extend to the cases where $m$ is polynomially or exponentially larger than $n$ [^5] Notice that our results for regret analysis and those in the adversarial setting hold regardless of the number of excess servers. In particular, as $m$ increases, the competitive ratio of our algorithm also increases correspondingly with a context-dependent increasing rate, and when $m=\omega(n)$, our algorithm no longer achieves a constant competitive ratio for the Euclidean space with $d\geq 3$. Moreover, relaxing our assumptions that servers outnumber requests and all requests must be served leads to an interesting variant where request rejection is permitted—a setting that bears similarities to the multi-unit prophet inequality problem [JMZ22]. Finally, it is appealing to relax the assumption that the cost function satisfies the triangle inequality, and we establish positive results for cost functions only approximately satisfying the triangle inequality in Appendix C.
@@ A.3 Proof of Lemma 4
Appendix B Postponed Proofs in Section 7.1
B.1 Proof of Lemma 12
Therefore,
which concludes that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing.
B.2 Proof of Lemma 13
concluding that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing.
Appendix C Stochastic Setting: Cost Functions Beyond Triangle Inequality
In previous sections, we consider metric space where the cost function satisfies the exact triangle inequality. In this section, we showcase the generality of our reduction for the stochastic setting by proving positive results when the cost function only approximately satisfies the triangle inequality. Specifically, we say that a cost function $\delta$ on $X$ satisfies the $\eta$-approximate triangle inequality for $\eta>0$ if for all $x,y,z\in X$,